Ed and Dave are going to divide a cash prize of $14,000 in a ratio of 2 to 5. How much will Dave receive?

Thanks.

Answer by stanbon(48558)

Ed and Dave are going to divide a cash prize of $14,000 in a ratio of 2 to 5. How much will Dave receive?
------------
2 to 5 is the same as 2x to 5x
---------
EQUATION:
2x+5x=14000
7x=14000
x=2000
---------
Dave has the larger amount so he has 5x=5*2000=$10000
===============
Cheers,
Stan H.